Waterway Corporation issued 342 shares of $10 par value common stock and 150 shares of $50 par value preferred stock for a lump sum of $19,656. The common stock has a market price of $20 per share, and the preferred stock has a market price of $100 per share.

Prepare the journal entry to record the issuance.
Answer #1

Journal entry

Date General Journal Debit Credit
Cash 19656
Common Stock (342*10) 3420
Paid in capital in excess of par-Common Stock (19656*6840/21840)-3420 2736
Preferred stock (150*50) 7500
Paid in capital in excess of par-Preferred stock 6000

Market value of Common Stock = 342*20 = 6840

Market value of preferred stock = 150*100 = 15000